2 raise to the fifth power times 2 raise to the third power = 2 raise to the __ power
Answer by lwsshak3(11628)

2 raise to the fifth power times 2 raise to the third power = 2 raise to the 8/15 power
2^(1/5)*2^(1/3)
add exponents
=2^(1/5+1/3)
=2^(3/15+5/15)
=2^(8/15)